OpenStudy (anonymous):

A solution of salt and water contains 75 grams of water per 150 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 30 milliliters of the solution. [Round off your answer to two decimal places]

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

75 grams of water is to 150 milliliters of the solution and X grams of water is to 30 milliliters of the solution by ratio and proportion 75/150=X/30 so you will get the value of X X=15 15 grams of water is to 30 milliliters of the solution by ratio and proportion again 1 mole of water is to 16 grams and X moles of water is to 15 grams which we solved earlier (30 milliliters) so 1/16=X/15 the X will be solved X=0.9375=0.94 moles

OpenStudy (anonymous):

here u go

OpenStudy (anonymous):

thankyou :)

OpenStudy (anonymous):

welcome